### v1

# Gravitational Waves from First-Order Phase Transitions Assisted by Temperature-Enhanced Scatterings

Abstract
Scatterings whose cross sections increase as the cosmic temperature decreases, known as temperature-enhanced scatterings, can have a significant impact on the thermal effective potential of scalar fields responsible for driving cosmological first-order phase transitions. In this work, we systematically investigate how the inclusion of temperature-dependent corrections to the effective potential affects key phase transition parameters, including the nucleation temperature, latent heat release, and inverse duration parameter. These modifications influence both the strength and duration of the phase transition, which in turn determine the properties of the resulting stochastic gravitational wave (GW) background. Employing semi-analytic computational methods, we evaluate the GW spectra generated under these conditions and compare our predictions with the projected sensitivities of forthcoming detectors such as LISA, DECIGO, and BBO. Our analysis demonstrates that temperature-enhanced scatterings can substantially strengthen the phase transition and produce GW signals that lie within the reach of future observational facilities. This study highlights the importance of temperature-dependent microphysical processes in shaping early Universe cosmological signatures.
1 Introduction
The origin of the baryon asymmetry of the Universe (BAU) remains one of the most compelling open questions in modern cosmology and particle physics [1, 2, 3, 4, 5, 6]. Numerous mechanisms have been proposed, including electroweak baryogenesis [7, 8, 9, 5, 10, 11, 12], leptogenesis [13, 14, 4, 15, 16, 17, 18], Affleck–Dine baryogenesis [19, 20, 21], and mechanisms involving Q-balls [22, 23, 24], but the precise dynamics that produced the observed asymmetry remain elusive.
A recent proposal [25] has highlighted an alternative possibility: baryogenesis through critical and super-critical scatterings, where the cross sections for interactions between non-relativistic particles and the thermal bath increase as the Universe cools. Such temperature-enhanced scatterings can sustain or even amplify asymmetry generation at late times, circumventing the limitations faced by conventional scenarios.
In parallel, first-order phase transitions (FOPTs) in the early Universe have attracted significant interest as potential sources of a stochastic gravitational wave (GW) background [26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. A strongly first-order phase transition proceeds through bubble nucleation, expansion, and collisions, together with plasma processes such as sound waves and turbulence [36, 37, 38, 39, 40, 41, 42, 43, 32, 44]. These mechanisms can generate GW signals within the reach of planned experiments such as LISA [45, 32, 46], DECIGO [47, 48], BBO [49], ET [50], and CE [51].
It is therefore natural to ask whether the same temperature-enhanced scatterings that can drive baryogenesis might also affect the dynamics of a FOPT, potentially strengthening the transition and enhancing its GW signal. Such scatterings modify the thermal effective potential of the scalar field driving the phase transition, thereby altering the critical and nucleation temperatures, the latent heat released, and the duration of the transition. These parameters directly control the amplitude and spectral shape of the resulting GW background [31, 32, 33, 52, 53, 10, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66].
In this work, we explore this connection in a general phenomenological framework. Rather than committing to a specific UV-complete particle physics model, we parametrize the impact of temperature-enhanced scatterings on the thermal effective potential by introducing an additional term,
$\Delta V(\phi,T)=c\,T^{p}\phi^{2},$ (1.1)
where $p<0$ captures the growing impact of scatterings as the temperature decreases. We analyze how this term shifts key phase transition parameters, namely the nucleation temperature $T_{n}$, the strength parameter $\alpha$, and the inverse duration parameter $\beta$ [31, 32, 52].
We then compute the resulting GW spectrum using standard semi-analytic techniques [36, 31, 32, 35, 43]. By scanning the parameter space $(c,p)$, we identify regions where the GW signal may be within reach of planned observatories such as LISA, DECIGO, and BBO.
Our results demonstrate that temperature-enhanced scatterings can significantly amplify a FOPT by increasing its strength parameter $\alpha$ and reducing its inverse duration parameter $\beta$, thereby generating an observable GW signal. This provides an intriguing complementarity: the same mechanism that could produce the BAU may also lead to testable predictions for GW experiments. This connection highlights the power of GW cosmology as a probe of physics beyond the Standard Model and motivates further exploration of explicit particle models that realize this scenario [63, 58, 57, 56, 59, 60, 61, 62, 66, 65, 64, 67, 68, 69, 70, 71, 72].
This paper is organized as follows. In Section 2, we present the theoretical framework and parameterization of temperature-enhanced scatterings and describe their effect on the thermal effective potential. Section 3 outlines the methodology used to compute the associated stochastic GW spectrum. In Section 4, we examine the parameter space and evaluate the GW signal for representative benchmark points. Our main results are presented and discussed in Section 5, followed by our summary and outlook in Section 6.
2 Framework and methodology
In this section, we describe the theoretical setup used to explore the impact of temperature-enhanced scatterings on cosmological first-order phase transitions and the resulting GW signals. Rather than constructing a specific ultraviolet-complete particle physics model, we adopt a phenomenological approach that captures the essential physics and allows us to scan systematically over the relevant parameter space.
2.1 Temperature-enhanced scatterings
Recently, it has been proposed that the baryon asymmetry of the Universe may originate from scatterings whose cross sections increase as the temperature of the Universe decreases [25]. Such scatterings can be described by a temperature-dependent cross section of the form
$\sigma(T)\propto T^{-n},$ (2.1)
with $n>0$, so that the cross section becomes larger at lower temperatures. This scaling may arise, for example, in models with light mediators, composite states, or nonperturbative effects that become important as the thermal bath cools.
The key insight of [25] is that these temperature-enhanced scatterings can sustain or even increase baryogenesis efficiency at late times, circumventing the usual washout and Boltzmann suppression that affect standard scenarios. We explore whether the same scatterings, through their impact on the thermal effective potential, can also modify the dynamics of a first-order phase transition, potentially leading to observable GW signals.
2.2 Parametrization of the thermal effective potential
To study the impact of temperature-enhanced scatterings on the scalar field $\phi$ responsible for the phase transition, we parametrize the thermal effective potential as
$V_{\mathrm{eff}}(\phi,T)=V_{0}(\phi)+\Delta V(\phi,T),$ (2.2)
where $V_{0}(\phi)$ is the zero-temperature potential and $\Delta V(\phi,T)$ encodes finite-temperature corrections.
The new ingredient is the contribution of temperature-enhanced scatterings, modeled by the additional term 1.1 with $c>0$ and $p<0$. The negative power $p$ captures the fact that scatterings become stronger as the temperature decreases, effectively increasing the thermal mass term for $\phi$ at lower temperatures. This phenomenological form allows us to explore a wide class of possible microscopic realizations in a model-independent way.
At high temperatures, this term is subdominant compared to the usual $a\,T^{2}\phi^{2}$ thermal corrections (with $a>0$), but as the Universe cools, the term $c\,T^{p}\phi^{2}$ can become significant, potentially delaying or strengthening the phase transition.
Figure 1 illustrates the evolution of the effective potential $V_{\mathrm{eff}}(\phi,T)$ at several representative temperatures. Solid lines include the temperature-enhanced scattering term, while dashed lines show the standard potential without this new contribution. As the temperature decreases, the presence of the temperature-enhanced scattering term deepens the broken phase minimum and can modify the critical temperature and nucleation temperature of the phase transition.
Figure 1: Evolution of the effective potential $V_{\mathrm{eff}}(\phi,T)$ as the Universe cools, shown at several temperatures $T$. Solid lines include the temperature-enhanced scattering term $cT^{p}\phi^{2}$, while dashed lines show the standard potential without this term. The new term deepens the broken phase minimum or delays the transition, illustrating how temperature-enhanced scatterings can impact the dynamics of the phase transition.
In Fig. 1, the temperatures $T=1.0,0.8,0.6$ are given in dimensionless units scaled by a reference temperature $T_{\mathrm{ref}}$, which can be interpreted as the characteristic energy scale of the model (for example, the zero-temperature vacuum expectation value of the scalar field or the critical temperature in the absence of the temperature-enhanced term). These values represent a snapshot of the cosmological evolution as the Universe cools down. At the highest temperature $T=1.0$, the effective potential favors the symmetric phase with the vacuum at $\phi=0$. As the temperature decreases to $T=0.8$, the broken minimum at $\phi\neq 0$ begins to form but may still be metastable. At the lowest temperature shown, $T=0.6$, the broken phase minimum becomes the global minimum, indicating that the phase transition has completed or is about to complete. The inclusion of the temperature-enhanced scattering term alters the depth and location of these minima, which can significantly affect the critical and nucleation temperatures as well as the strength of the phase transition. This modification of the effective potential will have a direct impact on the phase transition dynamics and GW signals, as discussed in subsequent sections.
2.3 Phase transition parameters
The dynamics of a cosmological first-order phase transition are typically characterized by several key parameters [31, 32, 33, 73, 74]:
  • Critical temperature $T_{c}$: the temperature at which the symmetric and broken minima of the thermal effective potential $V_{\mathrm{eff}}(\phi,T)$ become degenerate [75, 76, 9].
  • Nucleation temperature $T_{n}$: the temperature at which bubbles of the broken phase nucleate efficiently, defined by the condition $S_{3}(T_{n})/T_{n}\approx 140$, where $S_{3}(T)$ is the three-dimensional Euclidean bounce action [77, 78, 75]. This ensures at least one bubble per Hubble volume is nucleated.
  • Strength parameter $\alpha$: the ratio of the latent heat released during the phase transition to the radiation energy density at $T_{n}$, which controls the fraction of vacuum energy converted into bulk motion and GW [31, 52, 53].
  • Inverse duration parameter $\beta$: defined as $\beta/H\big{|}_{T_{n}}$, where $H$ is the Hubble parameter at $T_{n}$ [31, 32]. It measures how rapidly the transition completes; smaller $\beta/H$ corresponds to a slower, longer transition producing stronger GW signals.
In our framework, temperature-enhanced scatterings introduce an additional term of the form $c\,T^{p}\phi^{2}$ in the effective potential, with $p<0$ [79, 80, 74]. As the Universe cools, this term becomes increasingly significant, modifying the thermal evolution of $V_{\mathrm{eff}}(\phi,T)$ and directly affecting the bounce action $S_{3}(T)/T$, where $S_{3}(T)$ is the three dimensional Euclidean action. Details about the bounce action is given in Appendix A. As illustrated in Fig. 2, different values of $p$ change the temperature dependence of $S_{3}(T)/T$, shifting the nucleation temperature $T_{n}$, altering the duration of the phase transition (through $\beta/H$), and affecting its strength (through $\alpha$).
Figure 2: Temperature dependence of the bounce action $S_{3}(T)/T$ for different exponents $p$ in the temperature-enhanced scattering term. The horizontal dashed line at $S_{3}/T=140$ indicates the nucleation criterion. Red points mark the nucleation temperature $T_{n}$ where the bounce action drops below this threshold. The temperature on the x-axis is shown in units of the critical temperature, $T/T_{c}$, making the plot dimensionless. As $p$ varies, the dynamics of the phase transition change, shifting the nucleation temperature and affecting the strength and duration of the transition.
These phase transition parameters directly determine the amplitude, peak frequency, and shape of the stochastic GW background produced in the early Universe [31, 32, 74]. In the following sections, we numerically construct $V_{\mathrm{eff}}(\phi,T)$ including temperature-enhanced scatterings, extract these parameters, and explore their phenomenological consequences for GW signals.
2.4 Example UV Motivation
To illustrate how temperature-enhanced scatterings can arise from an explicit microphysical model, we sketch a minimal scenario involving a light mediator. Suppose the scalar field $\varphi$ responsible for the phase transition couples to a light scalar mediator $S$, which itself interacts with thermal bath fermions $\psi$. The relevant interaction Lagrangian is given by
$\mathcal{L}\supset-\frac{1}{2}m_{S}^{2}S^{2}-y_{S}\,S\,\bar{\psi}\psi-\mu\,S\,\varphi^{2}.$ (2.3)
In this setup, scatterings of the form $\psi\,\varphi\to\psi\,\varphi$ proceed through $t$-channel exchange of the mediator $S$. The corresponding amplitude contains the propagator term $1/(t-m_{S}^{2})$, which grows as the Mandelstam variable $t$ approaches zero at low temperatures. Consequently, the cross section behaves approximately as
$\sigma(\psi\,\varphi\to\psi\,\varphi)\sim\frac{y_{S}^{2}\,\mu^{2}}{(T^{2}+m_{S}^{2})^{2}},$ (2.4)
where the typical momentum transfer is set by the thermal energy, $|t|\sim T^{2}$. For temperatures $T\gg m_{S}$, this yields a scaling $\sigma(T)\propto T^{-4}$, providing a concrete realization of a cross section that grows as the Universe cools.
At finite temperature, such enhanced scatterings contribute an additional correction to the thermal effective potential for $\varphi$ through loop effects. Integrating out the thermal bath degrees of freedom generates an effective term of the form
$\Delta V(\varphi,T)\propto c\,T^{p}\,\varphi^{2},$ (2.5)
with $p<0$, consistent with our phenomenological parametrization in Eq. 1.1.
This minimal example demonstrates that temperature-enhanced scatterings can naturally emerge in models with light mediators and trilinear couplings. Such ingredients are commonplace in many dark sector or extended scalar sector scenarios, supporting the generality of our approach.
3 Gravitational wave prediction
3.1 GW sources from FOPTs
A FOPT in the early Universe generates a stochastic background of GWs through several distinct physical processes. The primary sources include bubble collisions, sound waves, and magnetohydrodynamic (MHD) turbulence, each contributing uniquely to the overall GW spectrum [31, 32, 73].
Bubble collisions arise from the dynamics of expanding true vacuum bubbles that nucleate and collide during the phase transition. These collisions produce GWs primarily through scalar field gradients localized at the bubble walls, as initially studied in Refs. [29, 81]. Although this mechanism can yield a distinct spectral shape, its contribution is often subdominant compared to fluid effects in many realistic scenarios.
The dominant source in most cosmological first-order phase transitions is the sound waves generated by the coherent bulk motion of the plasma. As expanding bubbles push the surrounding plasma, sound waves propagate and persist over timescales comparable to the Hubble time, efficiently sourcing gravitational radiation [37, 38]. The resulting GW signal from sound waves typically has a characteristic spectral peak and amplitude that depend sensitively on the strength and duration of the phase transition.
In addition, turbulent motion of the plasma and the associated magnetic fields induced after bubble collisions produce GWs through magnetohydrodynamic (MHD) turbulence [82, 83]. Although turbulence is expected to be a subdominant source compared to sound waves, it can contribute non-negligibly to the GW spectrum at high frequencies and alter the spectral shape.
The overall shape, peak frequency, and amplitude of the stochastic GW background generated by a first-order phase transition are controlled by the key parameters characterizing the transition: the strength parameter $\alpha$, which measures the ratio of vacuum energy released to the radiation energy density; the inverse duration parameter $\beta/H$, which quantifies the rapidity of the transition; the bubble wall velocity $v_{w}$; and the nucleation temperature $T_{n}$. In our framework, temperature-enhanced scatterings modify the thermal effective potential and thus affect the evolution of the bounce action. This leads to changes in these phase transition parameters, ultimately influencing the predicted GW signal.
In the following subsection, we detail the computational approach employed to quantitatively evaluate the contributions from these sources and to predict the resulting GW spectra.
3.2 Computation setup
To compute the stochastic GW spectrum generated by the first-order phase transition, we adopt the standard semi-analytic formalism developed in Refs. [31, 32, 36, 38]. In this framework, the total GW energy density spectrum today can be expressed as the sum of three main contributions:
$\Omega_{\mathrm{GW}}(f)h^{2}=\Omega_{\mathrm{col}}(f)h^{2}+\Omega_{\mathrm{sw}}(f)h^{2}+\Omega_{\mathrm{turb}}(f)h^{2},$ (3.1)
where the terms correspond respectively to GWs sourced by the collision of bubble walls ($\Omega_{\mathrm{col}}$), the long-lasting sound waves in the plasma ($\Omega_{\mathrm{sw}}$), and the turbulent motion of the plasma ($\Omega_{\mathrm{turb}}$) following the transition.
The shape and peak frequency of each component depend on the thermal parameters characterizing the phase transition. The peak frequency today can be approximately written as
$f_{\mathrm{peak}}\simeq 1.65\times 10^{-5}~{}\mathrm{Hz}\left(\frac{f_{*}}{\beta}\right)\left(\frac{\beta}{H}\right)\left(\frac{T_{n}}{100\,\mathrm{GeV}}\right)\left(\frac{g_{*}}{100}\right)^{1/6},$ (3.2)
where $T_{n}$ is the nucleation temperature, $\beta/H$ is the inverse duration parameter in units of the Hubble rate at $T_{n}$, $g_{*}$ is the number of relativistic degrees of freedom, and $f_{*}/\beta$ is an order-one numerical coefficient that depends on the detailed source dynamics.
The amplitude of each contribution is controlled by the strength parameter $\alpha$, the bubble wall velocity $v_{w}$, and the efficiency factors $\kappa_{i}(\alpha,v_{w})$ which quantify the fraction of the latent heat converted into bulk motion, sound waves, or turbulence. The bubble wall velocity $v_{w}$ is treated as a phenomenological parameter, typically varied within $v_{w}\in[0.6,1.0]$. The values of $\alpha$ and $\beta/H$ are extracted numerically from the temperature-dependent effective potential $V_{\mathrm{eff}}(\phi,T)$ described in earlier sections, which includes the temperature-enhanced scattering term $c\,T^{p}\phi^{2}$.
Our computation proceeds by first scanning over the parameters $p$ and $c$ to generate a set of effective potentials for different choices of the temperature-enhanced scattering term. For each parameter point, we compute the bounce action $S_{3}(T)/T$ as a function of temperature to extract the nucleation temperature $T_{n}$. Using $T_{n}$, we then determine the strength parameter $\alpha$ by calculating the released latent heat, and evaluate $\beta/H$ from the temperature derivative of the bounce action at nucleation:
$\beta/H=T_{n}\left.\frac{d(S_{3}/T)}{dT}\right|_{T=T_{n}}.$ (3.3)
We use the package cosmoTransitions, [84], to solve the above equation. Once these thermal parameters are known, individual contributions $\Omega_{\mathrm{col}}$, $\Omega_{\mathrm{sw}}$, and $\Omega_{\mathrm{turb}}$ to the GW spectrum are evaluated using semi-analytic formulae that encode the known dependence on $\alpha$, $\beta/H$, $v_{w}$, and the nucleation temperature $T_{n}$. This approach captures the essential physics of GW generation from expanding and colliding bubbles, as well as the subsequent sound waves and turbulence in the plasma.
Throughout this analysis, we systematically vary the exponent $p$ and the coefficient $c$ to explore how temperature-enhanced scatterings affect the key phase transition parameters and, consequently, the predicted GW spectra. The goal is to identify regions of parameter space where the resulting GW signals fall within the projected sensitivity ranges of future experiments such as LISA, DECIGO, and BBO.
3.3 Bubble wall friction estimate
The motion of expanding bubble walls during a first-order phase transition is subject to friction due to interactions with particles in the thermal plasma [53]. In our framework, the temperature-enhanced scatterings can modify this friction by increasing the effective interaction rate between the scalar field $\varphi$ and thermal bath particles.
At leading order, the friction coefficient $\eta$ is proportional to the thermal interaction rate,
$\eta\sim\sum_{i}\Delta m_{i}^{2}\,n_{i}(T)/\Gamma_{i},$ (3.4)
where $\Delta m_{i}^{2}$ is the change in the particle’s mass squared across the bubble wall, $n_{i}(T)$ is the thermal number density, and $\Gamma_{i}$ is the interaction rate. In the presence of enhanced scatterings, the rate $\Gamma_{i}$ increases due to the $T$-dependent cross section $\sigma(T)\propto T^{-n}$.
Consequently, the friction grows as the Universe cools. However, for the range of parameters we consider ($n\sim 1{-}4$), the friction enhancement remains moderate and does not prevent bubble expansion, provided the latent heat is sufficient to accelerate the wall.
To ensure that our semi-analytic treatment remains consistent, we have verified that the bubble wall velocity $v_{w}$ remains in the typical range $0.6\lesssim v_{w}\lesssim 1$, where sound wave production dominates the GW signal. A more detailed calculation of friction and wall dynamics would require a coupled treatment of Boltzmann transport equations and is beyond the scope of this study.
In the following section, we present the detailed numerical results, illustrating the dependence of the GW peak amplitude and frequency on the model parameters and discussing the phenomenological implications for experimental detection.
4 Parameter space analysis and phenomenology
In this section, we systematically explore the parameter space associated with the temperature-enhanced scattering term in the thermal effective potential,
$V_{\mathrm{eff}}(\phi,T)=V_{0}(\phi)+\Delta V_{\mathrm{thermal}}(\phi,T)+c\,T^{p}\phi^{2},$ (4.1)
where $c$ is a dimensionless coefficient controlling the strength of the scattering effect, and $p<0$ determines how this contribution scales with temperature. The new term captures scatterings whose cross sections increase as the Universe cools, enhancing the phase transition at late times. By scanning over the ranges
$p\in[-1.5,0],\quad c\in[0,c_{\max}],$ (4.2)
where $c_{\max}$ is chosen to preserve perturbativity and vacuum stability, we quantify how these parameters affect the dynamics of the electroweak-like phase transition and the resulting GW signatures.
Our analysis proceeds in three stages:
  • compute the temperature-dependent effective potential $V_{\mathrm{eff}}(\phi,T)$ across the parameter space,
  • extract the key phase transition parameters—critical temperature $T_{c}$, nucleation temperature $T_{n}$, strength parameter $\alpha$, and inverse duration parameter $\beta/H$,
  • predict the stochastic GW signal and assess its observability at upcoming experiments such as LISA, DECIGO, and BBO.
The results are summarized below.
4.1 Scanning the temperature-enhanced scattering parameters
For each pair $(p,c)$, we numerically compute $V_{\mathrm{eff}}(\phi,T)$ and track the evolution of its minima with temperature. The critical temperature $T_{c}$ is defined by the condition
$V_{\mathrm{eff}}(0,T_{c})=V_{\mathrm{eff}}(\phi_{\mathrm{min}}(T_{c}),T_{c}),$ (4.3)
where $\phi_{\mathrm{min}}(T)$ is the nontrivial minimum. The nucleation temperature $T_{n}$ is determined by solving
$\frac{S_{3}(T_{n})}{T_{n}}\approx 140,$ (4.4)
with $S_{3}(T)$ the three-dimensional Euclidean bounce action.
We also compute:
$\alpha$ $=\frac{\Delta\rho}{\rho_{\mathrm{rad}}(T_{n})},$ (4.5)
$\frac{\beta}{H}$ $=T_{n}\left.\frac{d}{dT}\left(\frac{S_{3}(T)}{T}\right)\right|_{T_{n}},$ (4.6)
where $\Delta\rho$ is the released latent heat and $\rho_{\mathrm{rad}}(T_{n})$ is the radiation energy density at $T_{n}$.
Figure 3 presents the results:
  • Left: The strength parameter $\alpha$ increases significantly as $p$ becomes more negative and $c$ reaches moderate values, reflecting enhanced symmetry breaking at lower temperatures.
  • Right: The inverse duration parameter $\beta/H$ correspondingly decreases, indicating slower transitions that last longer, which is favorable for GW production.
These trends show that the temperature-enhanced scattering term systematically enlarges the region of parameter space that supports strong, long-lasting first-order phase transitions.
Figure 3: Scan over the temperature-enhanced scattering parameters $p$ and $c$. Left: Strength parameter $\alpha$. Right: Inverse duration parameter $\beta/H$. Regions with larger $\alpha$ and smaller $\beta/H$ correspond to stronger and longer-lasting phase transitions, most pronounced for large negative $p$ and moderate $c$.
4.2 Impact on phase transition parameters
Beyond $\alpha$ and $\beta/H$, the temperature-enhanced scatterings also modify $T_{c}$ and $T_{n}$. Figure 4 shows how $T_{n}/T_{c}$ varies with $p$ for several benchmark values of $c$. As $p$ becomes more negative, $T_{n}$ shifts to lower values relative to $T_{c}$, since the scattering term becomes more important at cooler epochs, delaying bubble nucleation.
Figure 4: Nucleation temperature $T_{n}$ normalized to the critical temperature $T_{c}$ versus exponent $p$ for different values of $c$. Larger negative $p$ leads to later nucleation (lower $T_{n}/T_{c}$), as the scattering term increasingly favors the broken phase at lower temperatures.
We also observe the characteristic inverse correlation between $\alpha$ and $\beta/H$:
$\alpha\uparrow\quad\Longrightarrow\quad\beta/H\downarrow,$ (4.7)
illustrated in Figure 5. The temperature-enhanced scattering term shifts and broadens this correlation, populating the region with $\alpha\gtrsim 0.1$ and $\beta/H\lesssim 100$, where GW production is most efficient.
Figure 5: Correlation between the strength parameter $\alpha$ and inverse duration parameter $\beta/H$ across the scanned parameter space. Temperature-enhanced scatterings extend the region of strong and slow transitions (upper-left region), enhancing the predicted GW signal.
4.3 Gravitational wave signal and detector prospects
Using the extracted $\alpha$, $\beta/H$, and $T_{n}$, we compute the stochastic GW spectra following the standard formalism:
$\Omega_{\mathrm{GW}}h^{2}(f)\approx\Omega_{\mathrm{peak}}\,S(f/f_{\mathrm{peak}}),$ (4.8)
where $\Omega_{\mathrm{peak}}$ depends on $\alpha$, $\beta/H$, $T_{n}$, and other efficiency factors, and $S(f/f_{\mathrm{peak}})$ is a shape function determined by the physics of bubble collisions, sound waves, and turbulence.
Figure 6 shows the peak GW energy density $\Omega_{\mathrm{GW}}h^{2}$ versus $p$ for several values of $c$, compared against projected sensitivities of LISA, DECIGO, and BBO. We find that sufficiently large negative $p$ and moderate $c$ can generate GW signals exceeding the detectability thresholds of these observatories.
Figure 6: Peak GW energy density $\Omega_{\mathrm{GW}}h^{2}$ as a function of the exponent $p$ for various coefficients $c$. Shaded horizontal bands show the projected sensitivities of LISA, DECIGO, and BBO. Parameter choices with larger negative $p$ and moderate $c$ lead to detectable signals.
Finally, Figure 7 presents the full predicted GW spectra for representative benchmark points, demonstrating how variations in $(p,c)$ change the peak amplitude and frequency, and shift the spectra across the sensitivity windows of different experiments.
Figure 7: Heatmap of the predicted peak GW amplitude $\Omega_{\mathrm{GW}}h^{2}$ (arbitrary units) across the $(p,c)$ parameter space, where $p$ is the temperature exponent and $c$ the scattering coefficient in the temperature-enhanced term. The plot highlights how larger negative values of $p$ and moderate $c$ enhance the GW signal, potentially bringing it within the reach of future detectors such as LISA, DECIGO, and BBO. Contour lines illustrate regions with higher predicted amplitudes, emphasizing the phenomenologically promising parameter space.
Overall, our analysis shows that temperature-enhanced scatterings significantly broaden the viable parameter space for strong first-order phase transitions and observable GWs, providing a robust, testable signature for future GW observatories.
5 Results
In this section, we present the main findings of our analysis of temperature-enhanced scatterings during cosmological FOPTs. We show how these scatterings modify the phase transition strength, duration, and nucleation temperature, and analyze the consequences for the resulting stochastic GW spectra. Our results highlight the phenomenological potential of this mechanism for producing observable signals at upcoming GW observatories.
5.1 Impact on phase transition strength
We begin by quantifying the effect of the temperature-enhanced scattering term (1.1) on the phase transition parameters: the strength parameter $\alpha$, inverse duration parameter $\beta/H$, critical temperature $T_{c}$, and nucleation temperature $T_{n}$.
By scanning over a range of exponents $p\in[-1.5,0]$ and coefficients $c\in[0,1]$, we numerically evaluate the temperature-dependent effective potential and compute the bounce action $S_{3}(T)/T$ as described in Section 2. The nucleation temperature $T_{n}$ is determined by the condition (4.4) ensuring that at least one bubble nucleates per Hubble volume. The strength parameter, $\alpha$ (4.5) quantifies the fraction of latent heat released relative to the background radiation energy density, while the inverse duration parameter, $\frac{\beta}{H}$ (4.6) measures how rapidly the transition completes.
Our scans show several clear trends:
  • Larger negative values of $p$ shift the nucleation temperature $T_{n}$ to lower values relative to $T_{c}$, as the scattering term grows more significant at lower temperatures.
  • The strength parameter $\alpha$ increases for larger $|p|$ and moderate $c$, indicating stronger phase transitions with more latent heat release.
  • The inverse duration parameter $\beta/H$ tends to decrease in these regions, implying slower, longer-lasting transitions that are more efficient at generating GWs.
These results are visualized in the heatmaps shown earlier (Figs. 3, 4, and 5). The temperature-enhanced scattering term systematically extends the viable parameter space where the phase transition becomes strong and slow enough to produce potentially detectable GW signatures.
5.2 Gravitational wave spectra
Next, we translate these phase transition parameters into predictions for the stochastic GW background. The peak amplitude of the GW energy density spectrum can be estimated as (4.8), where $S(f/f_{\mathrm{peak}})$ is a spectral shape function, and $\Omega_{\mathrm{peak}}$ depends on $\alpha$, $\beta/H$, $T_{n}$, and other efficiency factors, following the approach described in Section 3.
The temperature-enhanced scatterings influence the GW spectrum in several key ways:
  • By lowering $T_{n}$, they shift the peak frequency $f_{\mathrm{peak}}\propto T_{n}$ to lower values, closer to the sensitive bands of space-based interferometers.
  • Stronger transitions (larger $\alpha$) increase the peak amplitude, making the signal easier to detect.
  • Smaller $\beta/H$ leads to slower phase transitions, enhancing the overall GW energy density.
Figures 6 and 7 illustrate these effects:
  • Figure 6 shows the peak GW amplitude $\Omega_{\mathrm{GW}}h^{2}$ as a function of the exponent $p$ for several values of $c$, overlaid with the projected sensitivities of LISA, DECIGO, and BBO. Regions with larger negative $p$ and moderate $c$ yield signals above the sensitivity thresholds.
  • Figure 7 displays the full predicted GW spectra for benchmark parameter choices, highlighting how the peak shifts in frequency and amplitude across the detector bands.
Taken together, these results demonstrate that temperature-enhanced scatterings can significantly reshape the GW spectrum produced during a first-order phase transition, and open up new parameter regions where the resulting signals may be within reach of next-generation GW detectors.
6 Discussion and Conclusion
In this work, we have examined the cosmological implications of temperature-enhanced scatterings on the dynamics of first-order phase transitions (FOPTs) in the early Universe. Motivated by recent observations that certain thermal scattering cross sections can increase as the Universe cools, we modeled this effect by introducing an additive term in the thermal effective potential of the form $c\,T^{p}\phi^{2}$ with $p<0$ and $c>0$. Such a term naturally becomes more significant at lower temperatures, reshaping the potential landscape and altering the properties of the phase transition.
Our analysis shows that temperature-enhanced scatterings can substantially affect the dynamics of a FOPT. As the temperature decreases, the additional term deepens the broken minimum and lowers the barrier between the symmetric and broken phases. This shifts the nucleation temperature $T_{n}$ downward relative to the critical temperature $T_{c}$, causing the transition to occur later, when the ambient radiation energy density is lower. The resulting increase in latent heat enhances the strength parameter $\alpha$. Simultaneously, the bounce action $S_{3}(T)/T$ is modified, which alters the tunneling rate and prolongs the transition duration. Consequently, the inverse duration parameter $\beta/H$ is reduced, producing a stronger and longer-lasting transition. Together, these effects favor the generation of a detectable stochastic gravitational wave (GW) background.
We carried out a numerical study scanning a broad region of the $(p,c)$ parameter space and computed the resulting GW spectra using standard semi-analytic treatments for sound wave and bubble collision sources. The amplitude and spectral features of the GW background are highly sensitive to both $\alpha$ and $\beta/H$. A larger $\alpha$ and smaller $\beta/H$ significantly boost the GW signal. Moreover, since the peak frequency depends on $T_{n}$, delayed transitions shift the spectrum to lower frequencies, often into the optimal sensitivity range of future space-based detectors such as LISA, DECIGO, and BBO.
Our parameter scans identify sizeable regions of the $(p,c)$ plane where GW signals from temperature-enhanced FOPTs would be within reach of these experiments. Benchmark points illustrate typical spectra with a broken power-law shape characteristic of acoustic and collision-induced sources. Notably, for $p\lesssim-1$ and moderate $c$, the impact can be large enough to distinguish such transitions from those expected in standard thermal scenarios. These results highlight the potential of GW measurements to probe temperature-dependent microphysics in the early Universe.
Beyond their phenomenological impact, temperature-enhanced scatterings can naturally arise in explicit particle physics models, for example involving new scalars, dark sectors, or extended gauge structures. The effective term $c\,T^{p}\phi^{2}$ could result from loop-level thermal corrections mediated by light hidden particles. Embedding this mechanism in concrete UV completions remains an important next step, enabling connections with other observables such as dark matter relic abundance, electroweak precision tests, and collider searches. Moreover, if a phase transition occurs near the MeV scale, the modified thermal history could affect Big Bang Nucleosynthesis (BBN) and the Cosmic Microwave Background (CMB), motivating a more detailed study of cosmological consistency.
Our findings also motivate refinements of the theoretical treatment. These include accounting for higher-loop corrections to the thermal potential, incorporating plasma effects in modeling the bubble wall velocity, and performing non-perturbative lattice simulations to validate the tunneling dynamics in the presence of non-trivial temperature dependence. Such improvements will be vital for making robust predictions for GW observables.
We note that temperature-enhanced scatterings that amplify the phase transition also increase the friction acting on expanding bubble walls. As discussed in Section 3, we have verified that this increased interaction rate moderately raises the friction but does not prevent bubble expansion within the parameter space considered. The bubble wall velocity remains in the regime where sound waves dominate the GW production, ensuring that our semi-analytic predictions are reliable. A dedicated transport analysis could refine these estimates and will be an interesting direction for future study.
We also briefly comment on cosmological consistency. In our scenario, the nucleation temperature $T_{n}$ typically remains well above the MeV scale, ensuring that the phase transition completes safely before the onset of BBN. The modified thermal history thus does not disrupt standard light element abundances. Furthermore, the temperature-enhanced scatterings do not inject significant additional entropy or dark radiation at late times, preserving consistency with CMB bounds on the effective number of relativistic species $N_{\mathrm{eff}}$. A more detailed investigation of possible late-time signatures will be an important follow-up.
In conclusion, temperature-enhanced scatterings represent a simple yet powerful modification of the thermal scalar potential, capable of generating strong first-order phase transitions with observable GW signatures. This mechanism broadens the theoretical landscape of viable early-Universe scenarios and provides promising targets for upcoming GW observatories. A future detection of a stochastic GW background consistent with our predictions would not only open a new window on the microphysics of the early Universe but also reveal thermal processes beyond the reach of conventional probes, offering a unique complement to laboratory experiments.
7 Acknowledgment
The work of AC was supported by the Japan Society for the Promotion of Science (JSPS) as a part of the JSPS Postdoctoral Program (Standard), grant number JP23KF0289.
Appendix A Details of bounce action calculation
In this appendix, we provide a detailed account of the methodology used to compute the three-dimensional Euclidean bounce action $S_{3}(T)$, which governs the dynamics of bubble nucleation during a cosmological first-order phase transition. This quantity plays a central role in determining the nucleation temperature $T_{n}$, as the probability per unit volume per unit time of nucleating a critical bubble is exponentially suppressed by the factor $e^{-S_{3}(T)/T}$. Specifically, the nucleation rate is approximated by
$\Gamma(T)\simeq T^{4}\left(\frac{S_{3}(T)}{2\pi T}\right)^{3/2}e^{-S_{3}(T)/T}.$ (A.1)
The nucleation temperature $T_{n}$ is typically defined as the temperature at which $S_{3}(T)/T\approx 140$, ensuring that on average at least one bubble nucleates within a Hubble volume.
To compute $S_{3}(T)$, we solve for the so-called bounce solution $\phi(r)$, a spherically symmetric solution of the equation of motion derived from the effective potential $V_{\mathrm{eff}}(\phi,T)$. This equation reads
$\frac{d^{2}\phi}{dr^{2}}+\frac{2}{r}\frac{d\phi}{dr}=\frac{dV_{\mathrm{eff}}(\phi,T)}{d\phi},$ (A.2)
with boundary conditions
$\lim_{r\to\infty}\phi(r)=0,\quad\left.\frac{d\phi}{dr}\right|_{r=0}=0,$ (A.3)
where $r$ denotes the radial coordinate in three-dimensional Euclidean space.
The effective potential $V_{\mathrm{eff}}(\phi,T)$ in our scenario includes the temperature-enhanced scattering term of the form $c\,T^{p}\phi^{2}$. This term is particularly significant because for negative exponents $p<0$, it becomes more pronounced at lower temperatures, effectively deepening the broken minimum and lowering the barrier between the symmetric and broken phases as the Universe cools.
We solve the bounce equation numerically using the overshoot/undershoot method. This procedure involves guessing an initial value $\phi(0)$ at the center of the bubble and integrating outward. If the field overshoots the false vacuum at large $r$, the initial value is decreased; if it undershoots, the initial value is increased. Through iterative refinement, this method converges to the solution that asymptotically approaches the false vacuum. Once the bounce profile $\phi(r)$ is obtained, the action is computed as
[…]
Figure 8 illustrates representative bounce profiles $\phi(r)$ computed at a fixed temperature for several values of the exponent $p$. We observe that larger negative values of $p$ lead to sharper and more localized profiles. This reflects the fact that the temperature-enhanced scattering term deepens the broken minimum at low temperatures, which in turn lowers the bounce action $S_{3}(T)$ and facilitates earlier nucleation.
Figure 8: Example bounce solutions $\phi(r)$ at a fixed temperature, computed for different values of the exponent $p$ in the temperature-enhanced scattering term. More negative $p$ values result in a deeper broken minimum and a sharper, more localized field profile, highlighting how temperature-enhanced scatterings alter the nucleation dynamics.
Throughout our numerical implementation, we carefully discretize the radial coordinate to ensure that the field settles to the false vacuum at large $r$, and check that the computed action $S_{3}(T)$ is stable under changes of the grid spacing and boundary size. The potential $V_{\mathrm{eff}}(\phi,T)$ is evaluated at each temperature, allowing us to track the evolution of the bounce action across the relevant temperature range.
These calculations, grounded in the semiclassical formalism originally developed by Coleman and Linde [85, 86], are central to determining the key phase transition parameters, including the nucleation temperature $T_{n}$, the strength parameter $\alpha$, and the inverse duration parameter $\beta/H$. These, in turn, form the basis for predicting the stochastic GW background, which is a central focus of this work.